Tohban Report 2014-03-05

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 19 Feb 2014
End Date: 26 Feb 2014
Tohban: Nicole Duncan
Tohban email:
Next Tohban: Mitsuo Oka
List all reports


Solar Activity

Solar activity increased through the week, especially in the past few days where we have many C-class events. There are currently 10 ARs on the disk, primarily populating the western disk. The most active region is 11991, just west of disk center. This region has a ~10% chance of an X-class event and is currently in a beta-gamma-delta configuration. Stereo-B shows 1 or 2 ARs that will rotate onto the disk later next week.

How many GOES flares occurred?

 Flares above B, C, M, X class were      0    31     3     1

And how many of these are listed in the RHESSI flare list?

 Flares above B, C, M, X class were      0    16     1     1

There were RHESSI flares/GOES flares 379 / 51 over the time range 25-Feb-14 04-Mar-14

Memory Management

The increased solar activity filled the SSR to a maximum ~30% over the weekend. During this time we emptied to ~12% at the end of the passes. We changed decimation to ACTIVE/VIGOROUS on Monday and added a few more passes to reduce the fill level. We're now at ~3% SSR. We are still taking rear events at night, with no fronts.

Spacecraft Status

The cold plate 1 and 2 are at 118 and 117K, putting them in the yellow zone.

Data Gaps


Detector issues

-HV commanding on D9r: The experimental decrease of D9 voltage yielded confusing results. The voltage reduction surprisingly increased the reset rates ( ~6500 to ~9000) in D9 without any noticeable change to D8. There was no issue when we lowered/raised HV on D9, and the reset rates returned to ~6500. It was decided to simply raise the thresholds on D6f, D8r and D9r.

-D9r threshold change: This also yielded confusing results when D9 rear fast valids shot up (10,000 --> 50,000) as a result of the increased threshold. Over the intervening week, the D9 rear fast valids have slowly come back down to ~5,000 with no additional commanding.

-D8r threshold change: fast valids dramatically decreased from 100,000 --> 1,000 counts. Within the past week its drifted back up to ~5,000 counts.

-D6f threshold change: fast rates dropped from 50,000 --> 5,000 due to the threshold change. But, in the past week its drifted back up to ~10,000 counts.

-D4 & D6 rear fast valids: No commanding was done to these segments, but their rear fast valids have risen to to ~1,000 --> 5,000 counts over the week, which is now comparable to D9r fast rates. Their lifetimes are ~93% and 98% respectively. These two detectors should be monitored by the next tohban.

-D9r spike: Feb 27, 12:21 UT. This was before the threshold commanding.

-D2 continues to show signs of segmentation, with rear events only dropping out during SAA and active rear reset rates at other times.

- To do this week: D6f commanding will be done ~Friday.

Spacecraft Management

2014/064: Battery 1 is at 80%, in the yellow.

2014/058 19:25 UTC: Threshold changes G6 front fast: 0x70 to 0x80 G8 rear fast: 0x45 to 0x60 G9 rear fast: 0x80 to 0x90

2014/062 21:51 UTC. Decimation change to Acive/Vigorous

Decimation Active/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Disabled (4 minutes before and after eclipse)
Night time data (rears) Enabled
Require extra passes? Yes
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? D4, D6, D8, D9 elevated rear count rates
Personal tools